f(x) = x^2 + 3 , g(x) = square root x-2

Find f(g(x)).
1
Expert's answer
2014-05-10T08:17:48-0400
f(g(x))=(g(x))^2+3=((x-1)^(1/2))^2+3=|x|+3.